# Spatio-temporal methods for estimating subsurface ocean thermal response to tropical cyclones

### v3

@@ 4.2.1 Pairing process
We follow closely the pairing process introduced in Cheng, Zhu and Sriver (2015). Recall that the tropical cyclone tracks are available at six-hourly time resolution and can be made continuous in time and space through straight line interpolation. Along each of these TC tracks, we seek Argo profiles that reported within $\pm 8^{\circ}$ (longitude/latitude) and within $12$ days before to $20$ days after each location on the tropical cyclone track. We partition these profiles into the baseline set (12-12 to 2-2 days relative to the TC) and the signal set (2-2 to +20+20 days relative to the TC). These two sets contain the profiles that are incidental to the TC. The baseline and signal time ranges are chosen primarily to accommodate sampling constraints posed by the Argo network; specifically, most floats sample on a ten-day frequency. Therefore, the baseline time range of 12-12 to 2-2 days ensures that each float can contribute at least one baseline profile. The signal time range allows for at least two signal profiles from a given float (assuming they fit a nearness condition detailed in the sequel). The signal profiles begin at 2-2 days to hedge against storm effects prior to passage and is consistent with the pairing process introduced by Cheng, Zhu and Sriver (2015).
@@
We impose an additional technical restriction with respect to the lineage of a baseline profile, which is defined to be the set of all signal profiles associated with a distinct baseline profile and that baseline profile itself. We require that all profiles in the lineage of a baseline profile must be separated in time by at least 72 hours, sparsifying the lineage. To our knowledge, Cheng, Zhu and Sriver (2015) imposes no such constraint. The choice of this separation is for a particular reason: we know that pairs of profiles close in time tend to be highly correlated, but our Gaussian process model (Kuusela and Stein, 2018) for estimating the between-pair correlations (a prerequisite for properly handling them in the final thin plate spline fit) may not provide reliable estimates in the rare cases where profiles occur more frequently than three days apart. Following a filter on minimum time separation within each lineage, we retain 16025 profile pairs across all five ocean regions: 2757 in the North Atlantic, 2897 in the East Pacific, 778 in the Indian Ocean, 2185 in the Southern Hemisphere, and 7409 in the West Pacific.
@@ 4.3 Seasonal mean field estimation
Although the temperature differences induced by tropical cyclone passage occur on relatively small temporal scales, there is a nontrivial seasonal effect experienced by the ocean during the tropical cyclone season, at the time scales considered in our analysis. We focus here on the methodology for estimating the seasonal cycle and fully detail the resulting empirical findings in Section 5. In order to properly account for the seasonal effect, we adapt previous methodology (Ridgway, Dunn and Wilkin, 2002; Roemmich and Gilson, 2009) to estimate a global mean field specifically for the “counterfactual,” i.e., we characterize the state of ocean temperatures during the TC season, using the non-TC profiles of Section 4.1.2.
Using the non-TC profiles of Section 4.1.2, we estimate, separately for each pressure level $z=10,20,\dotsc,200$ and for the vertically averaged temperature values, a global mean field through a variant of the Roemmich–Gilson climatology. The Roemmich–Gilson climatology (Roemmich and Gilson, 2009) was among the first to take advantage of the spatial coverage of the Argo array to characterize the mean state and annual cycle of temperature in the global upper ocean. Specifically, they used a local linear regression model (Ridgway, Dunn and Wilkin, 2002) for estimating the climatological mean temperature as a function of location $(x_{\mathrm{lon}},x_{\mathrm{lat}})$, pressure level $z$, and day of year $t$. This model is written down explicitly in Kuusela and Stein (2018), and our model, which differs marginally from that of Roemmich and Gilson (2009) in that we do not weigh observations based on distance from the center point $(x_{\mathrm{lon}}^{*},x_{\mathrm{lat}}^{*})$, nor do we use data from pressure levels $z^{\prime}\neq z$ to form estimates at pressure level $z$, is elucidated in Equation (6).
@@
and formally described, in terms of the functional ANOVA model, as
@@
(a) $m_{{\boldsymbol{x}}^{\prime}}({\boldsymbol{x}}^{\prime},t_{\text{signal{}}})-m_{{\boldsymbol{x}}^{\prime}}({\boldsymbol{x}^{\prime}},t_{\text{baseline{}}})$
(b) Fitted $\phi$ parameters
@@ 4.5 Thin plate splines
Having learned a model for the $(a(\boldsymbol{x},t_{\text{signal{}}})-a(\boldsymbol{x},t_{\text{baseline{}}}))$ term of (11) through the Gaussian process model of Section 4.4, we find ourselves properly positioned to estimate the TC-induced signal $s(d,\tau)$. This we do with a thin plate spline, a flexible nonparametric technique for fitting a surface in two dimensions, subject to a curvature penalty and observation weights. Specifically, we smooth the seasonally adjusted temperature differences using a fixed-knot thin plate spline smoother (Duchon, 1977; Wahba, 1980, 1990; Green and Silverman, 1994; Nychka, 2000; Wood, 2017), in which we use the covariances estimated in Section 4.4 to account for heteroskedasticity and cross-observation dependence (Rao, 1973; Draper and Smith, 1998; Nychka, 2000; Ruppert, Wand and Carroll, 2003). As with the local mean field model of Section 4.3 and the Gaussian process model of Section 4.4, this procedure is conducted separately at each pressure level $z\in\{10,20,\dotsc,200\}$ and for the vertically averaged profiles.
@@
The first term in (22) is a generalized least squares (Rao, 1973; Draper and Smith, 1998; Ruppert, Wand and Carroll, 2003) data-fit criterion, and the second term is a roughness penalty (Meinguet, 1979; Green and Silverman, 1994; Wood, 2017), with $\tilde{\mathbf{R}},\mathbf{L},{\mathbf{S}}$ defined in the sequel. The parameter $\lambda>0$ mediates the competing demands of the data fidelity and roughness penalty terms.
@@ 4.5.1 Construction of the smoother
where $\hat{\boldsymbol{\delta}}$ are the coefficients corresponding to the radial basis functions centered at the fixed knots and $\hat{\boldsymbol{\beta}}$ are the coefficients corresponding to the constant and linear terms. Predictions at a new set of locations $\hat{\boldsymbol{\xi}}_{1},\dotsc,\hat{\boldsymbol{\xi}}_{N}\in\mathbb{R}^{2}$ may be obtained as follows:
$\hat{\boldsymbol{y}}=\begin{bmatrix}\hat{\mathbf{R}}&\hat{\mathbf{L}}^{\top}\end{bmatrix}\begin{bmatrix}\hat{\boldsymbol{\delta}}\\ \hat{\boldsymbol{\beta}}\end{bmatrix}$ (27)
@@
4.5.2 Prediction variances and confidence intervals
Prediction variances and confidence intervals follow for the thin plate spline estimates through the linearity of the smoother and the assumption of Gaussian noise. Denote by ${\boldsymbol{\Sigma}}\in\mathbb{R}^{n\times n}$ the observation covariance matrix $\operatorname{\mathrm{cov}}(\boldsymbol{y})$ from (21). The covariance matrix of the predictions is given by:
$\operatorname{\mathrm{cov}}(\hat{\boldsymbol{y}})=\begin{bmatrix}\hat{\mathbf{R}}&\hat{\mathbf{L}}^{\top}\end{bmatrix}\mathbf{A}^{-1}\begin{bmatrix}\tilde{\mathbf{R}}^{\top}\mathbf{W}\\ \mathbf{L}\mathbf{W}\\ \end{bmatrix}{\boldsymbol{\Sigma}}\begin{bmatrix}\tilde{\mathbf{R}}^{\top}\mathbf{W}\\ \mathbf{L}\mathbf{W}\\ \end{bmatrix}^{\top}\mathbf{A}^{-1}\begin{bmatrix}\hat{\mathbf{R}}&\hat{\mathbf{L}}^{\top}\end{bmatrix}^{\top}$ (30)
@@ 5.1 Seasonal mean field adjustment
The warming effect in the seasonal mean field difference is a consistent feature in the pressure levels $z=30$ through $z=150$. The pressure levels $z\in\{10,20\}$ do not show a warming pattern and might even exhibit a slight cooling effect, while the pressure levels $z\geq 160$ show only a small contribution of the seasonal signal to differences in paired profiles. These phenomena are consistent with the spatial structure of the seasonal cycle (Roemmich and Gilson, 2009). An adjustment for seasonal warming was not performed in a previous analysis of ocean thermal response using paired Argo profiles (Cheng, Zhu and Sriver, 2015); as a result, some of the warming effects previously observed over longer time spans may in fact have been due to seasonal variation. The full sequence of results at all twenty pressure levels is deferred to Section 1 of the supplementary material (Hu et al., 2020).
@@ 5.2 Ocean thermal response as a continuous function of time and cross-track distance
Finally, in Figure 7, we perform a pointwise level $\alpha=0.05$ hypothesis test on the thin plate spline fit, using estimates of prediction variance derived in Section 4.5. The issues of performing a pointwise hypothesis test with $N=40000$ points are manifold, and we defer discussion of extensions through which we could make the significance test more sophisticated to Section 6. Nonetheless, we see in Figure 7 a simple fact: that the predicted cooling effect induced by tropical cyclones is of a greater magnitude than the variability of the predictions. Moreover, this cooling effect persists for a sustained period of time, stretching into two weeks following the tropical cyclone passage. We also see a rightward bias in the cooling effect, which may be attributed to the storm’s direction of rotation. [^2] Due to the Coriolis effect, TCs in the Northern Hemisphere will induce a greater cooling effect on the right side of the cross-track angle parameterization than on the left side, due to the combined effects of forward movement and counterclockwise wind rotation. TCs in the Southern Hemisphere, which experience clockwise wind rotation, will induce a greater cooling effect on the left side. In order to display this effect in a consistent manner across hemispheres, we flip the sign of the cross-track angle for TCs in the Southern Hemisphere when producing Figures 612.
@@
This phenomenon may be observed in Figure 8. In Figure 8 we recall the smoothed temperature differences at depth $z=10$, in which a dramatic cooling effect is depicted. At the deeper pressure level of $z=60$, however, we see a mixture of both cooling and warming; namely, a net increase in temperature is observed at $d\in(2,4)$, $\tau\in(0,6)$. Finally, we see in Figure 8 that by $z=150$, the warming effect has disappeared, while the cooling effect remains. We note that the characterization of this warming effect as a continuous function of both cross-track angle and time since TC passage is made possible by our methodological framework and was not possible in previous literature (Cheng, Zhu and Sriver, 2015).
The discussion of vertical mixing is further developed in Sections 5.3, 5.4, and $5.$ The full set of covariance-reweighed thin plate spline fits, at all pressure levels $z=10,\dotsc,200$, is deferred to Section 2 of the supplementary material (Hu et al., 2020).
@@ 5.4 Depth-cross-track parameterization
We further demonstrate that our continuous-time methodology captures the temporally binned results of Cheng, Zhu and Sriver (2015), whose profile pairing process we used in this paper. To introduce temporal binning, we take the covariance-reweighed thin plate spline estimates from Section 5.2 across pressure levels $z=10,\dotsc,200$ and split them into two bins: $\tau\in[0,+3)$ and $\tau\in[+3,+20)$, termed the “forced” stage and “recovery” stage by Cheng, Zhu and Sriver (2015). For each of the two bins, we perform the same averaging process as in Section 5.3 to marginalize the estimates over the $\tau$ axis and obtain $\alpha=0.05$ pointwise significance tests.
@@
Several noteworthy phenomena are present in Figure 10. In Figure 10, we observe both the primary cooling effect and the secondary warming effect in the subsurface on the right side of the storm due to vertical mixing, as discussed in Sections 5.2 and 5.3. We also observe a sustained cooling effect in Figure 10, during the period called the “recovery” stage in Cheng, Zhu and Sriver (2015). While the primary cooling effect observed in the two time periods is qualitatively consistent with the results presented by Cheng, Zhu and Sriver (2015), their results also show a strong warming effect on both sides of the storm during the recovery stage (Figure 10 in Cheng, Zhu and Sriver, 2015), which is absent from our estimate. Overall, it appears that the recovery stage estimate in Cheng, Zhu and Sriver (2015) is “warm biased” in comparison to our estimate. We suspect that this difference is due to our methodology’s more careful accounting of the seasonal mean shift as described in Section 4.3.
@@ 5.5 Isosurfaces of the thin plate spline estimates
(a) $(\theta,\varphi)=(21.0^{\circ},144.1^{\circ})$
(b) $(\theta,\varphi)=(25.1^{\circ},16.1^{\circ})$
(c) $(\theta,\varphi)=(18.6^{\circ},176.4^{\circ})$
@@
For each set of isosurfaces, we first mask the thin plate spline estimates that fail to reject a pointwise level $\alpha=0.05$ hypothesis test, using the prediction variances from Section 4.5.2. We then apply the marching cubes algorithm (Lorensen and Cline, 1987; Lewiner et al., 2012; Virtanen et al., 2020) to identify a positive and a negative isosurface of magnitude $0.3^{\circ}\mathrm{C}$. The positive isosurface is colored in red, and the negative in blue.

### v4

@@ 3.2.1 Pairing process
We follow closely the pairing process introduced in Cheng et al. (2015). Recall that the tropical cyclone tracks are available at six-hourly time resolution and can be made continuous in time and space through straight line interpolation. Along each of these TC tracks, we seek Argo profiles that reported within $\pm 8^{\circ}$ (longitude/latitude) and within $12$ days before to $20$ days after each location on the tropical cyclone track. We partition these profiles into the baseline set (12-12 to 2-2 days relative to the TC) and the signal set (2-2 to +20+20 days relative to the TC). These two sets contain the profiles that are incidental to the TC. The baseline and signal time ranges are chosen primarily to accommodate sampling constraints posed by the Argo network; specifically, most floats sample on a ten-day frequency. Therefore, the baseline time range of 12-12 to 2-2 days ensures that each float can contribute at least one baseline profile. The signal time range allows for at least two signal profiles from a given float (assuming they fit a nearness condition detailed in the sequel). The signal profiles begin at 2-2 days to hedge against storm effects prior to passage and is consistent with the pairing process introduced by Cheng et al. (2015).
@@
We impose an additional technical restriction with respect to the lineage of a baseline profile, which is defined to be the set of all signal profiles associated with a distinct baseline profile and that baseline profile itself. We require that all profiles in the lineage of a baseline profile must be separated in time by at least 72 hours, sparsifying the lineage. To our knowledge, Cheng et al. (2015) imposes no such constraint. The choice of this separation is for a particular reason: we know that pairs of profiles close in time tend to be highly correlated, but our Gaussian process model (Kuusela and Stein, 2018) for estimating the between-pair correlations (a prerequisite for properly handling them in the final thin plate spline fit) may not provide reliable estimates in the rare cases where profiles occur more frequently than three days apart. Following a filter on minimum time separation within each lineage, we retain 16025 profile pairs across all five ocean regions: 2757 in the North Atlantic, 2897 in the East Pacific, 778 in the Indian Ocean, 2185 in the Southern Hemisphere, and 7409 in the West Pacific.
@@ 3.3 Seasonal mean field estimation
Although the temperature differences induced by tropical cyclone passage occur on relatively small temporal scales, there is a nontrivial seasonal effect experienced by the ocean during the tropical cyclone season, at the time scales considered in our analysis. We focus here on the methodology for estimating the seasonal cycle and fully detail the resulting empirical findings in Section 4. In order to properly account for the seasonal effect, we adapt previous methodology (Ridgway et al., 2002; Roemmich and Gilson, 2009) to estimate a global mean field specifically for the “counterfactual,” i.e., we characterize the state of ocean temperatures during the TC season, using the non-TC profiles of Section 3.1.2.
Using the non-TC profiles of Section 3.1.2, we estimate, separately for each pressure level $z=10,20,\dotsc,200$ and for the vertically averaged temperature values, a global mean field through a variant of the Roemmich–Gilson climatology. The Roemmich–Gilson climatology (Roemmich and Gilson, 2009) was among the first to take advantage of the spatial coverage of the Argo array to characterize the mean state and annual cycle of temperature in the global upper ocean. Specifically, they used a local linear regression model (Ridgway et al., 2002) for estimating the climatological mean temperature as a function of location $(x_{\mathrm{lon}},x_{\mathrm{lat}})$, pressure level $z$, and day of year $t$. This model is written down explicitly in Kuusela and Stein (2018), and our model, which differs marginally from that of Roemmich and Gilson (2009) in that we do not weigh observations based on distance from the center point $(x_{\mathrm{lon}}^{*},x_{\mathrm{lat}}^{*})$, nor do we use data from pressure levels $z^{\prime}\neq z$ to form estimates at pressure level $z$, is elucidated in Equation (6).
@@
and formally described, in terms of the ANOVA decomposition, as
@@
((a)) $m_{{\boldsymbol{x}}^{\prime}}({\boldsymbol{x}}^{\prime},t_{\text{signal{}}})-m_{{\boldsymbol{x}}^{\prime}}({\boldsymbol{x}^{\prime}},t_{\text{baseline{}}})$
((b)) Fitted $\phi$ parameters
@@ 3.5 Thin plate splines
Having learned a model for the $(a(\boldsymbol{x},t_{\text{signal{}}})-a(\boldsymbol{x},t_{\text{baseline{}}}))$ term of (11) through the Gaussian process model of Section 3.4, we find ourselves properly positioned to estimate the TC-induced signal $s(d,\tau)$. This we do with a thin plate spline, a flexible nonparametric technique for fitting a surface in two dimensions, subject to a curvature penalty and observation weights. Specifically, we smooth the seasonally adjusted temperature differences using a fixed-knot thin plate spline smoother (Duchon, 1977; Wahba, 1980, 1990; Green and Silverman, 1994; Nychka, 2000; Wood, 2017), in which we use the covariances estimated in Section 3.4 to account for heteroskedasticity and cross-observation dependence (Rao, 1973; Draper and Smith, 1998; Nychka, 2000; Ruppert et al., 2003). As with the local mean field model of Section 3.3 and the Gaussian process model of Section 3.4, this procedure is conducted separately at each pressure level $z\in\{10,20,\dotsc,200\}$ and for the vertically averaged profiles.
@@
The first term in (22) is a generalized least squares (Rao, 1973; Draper and Smith, 1998; Ruppert et al., 2003) data-fit criterion, and the second term is a roughness penalty (Meinguet, 1979; Green and Silverman, 1994; Wood, 2017), with $\tilde{\mathbf{R}},\mathbf{L},{\mathbf{S}}$ defined in the sequel. The parameter $\lambda>0$ mediates the competing demands of the data fidelity and roughness penalty terms.
@@ 3.5.1 Construction of the smoother
where $\hat{\boldsymbol{\delta}}$ are the coefficients corresponding to the radial basis functions centered at the fixed knots and $\hat{\boldsymbol{\beta}}$ are the coefficients corresponding to the constant and linear terms. Estimates of the signal $s(d,\tau)$ at a new set of locations $\hat{\boldsymbol{\xi}}_{1},\dotsc,\hat{\boldsymbol{\xi}}_{N}\in\mathbb{R}^{2}$ may be obtained as follows:
$\hat{\boldsymbol{s}}=\begin{bmatrix}\hat{\mathbf{R}}&\hat{\mathbf{L}}^{\top}\end{bmatrix}\begin{bmatrix}\hat{\boldsymbol{\delta}}\\ \hat{\boldsymbol{\beta}}\end{bmatrix}$ (27)
@@
3.5.2 Uncertainty quantification
Estimate variances and confidence intervals follow for the thin plate spline estimates through the linearity of the smoother and the assumption of Gaussian noise. Denote by ${\boldsymbol{\Sigma}}\in\mathbb{R}^{n\times n}$ the observation covariance matrix $\operatorname{\mathrm{cov}}(\boldsymbol{y})$ from (21). The covariance matrix of the estimates is given by:
$\operatorname{\mathrm{cov}}(\hat{\boldsymbol{s}})=\begin{bmatrix}\hat{\mathbf{R}}&\hat{\mathbf{L}}^{\top}\end{bmatrix}\mathbf{A}^{-1}\begin{bmatrix}\tilde{\mathbf{R}}^{\top}\mathbf{W}\\ \mathbf{L}\mathbf{W}\\ \end{bmatrix}{\boldsymbol{\Sigma}}\begin{bmatrix}\tilde{\mathbf{R}}^{\top}\mathbf{W}\\ \mathbf{L}\mathbf{W}\\ \end{bmatrix}^{\top}\mathbf{A}^{-1}\begin{bmatrix}\hat{\mathbf{R}}&\hat{\mathbf{L}}^{\top}\end{bmatrix}^{\top}$ (30)
@@ 4.1 Seasonal mean field adjustment
The warming effect in the seasonal mean field difference is a consistent feature in the pressure levels $z=30$ through $z=150$. The pressure levels $z\in\{10,20\}$ do not show a warming pattern and might even exhibit a slight cooling effect, while the pressure levels $z\geq 160$ show only a small contribution of the seasonal signal to differences in paired profiles. These phenomena are consistent with the spatial structure of the seasonal cycle (Roemmich and Gilson, 2009). An adjustment for seasonal warming was not performed in a previous analysis of ocean thermal response using paired Argo profiles (Cheng et al., 2015); as a result, some of the warming effects previously observed over longer time spans may in fact have been due to seasonal variation. The full sequence of results at all twenty pressure levels is deferred to Section 1 of the supplementary material (Hu et al., 2020).
@@ 4.2 Ocean thermal response as a continuous function of time and cross-track distance
Finally, in Figure 7, we perform a pointwise level $\alpha=0.05$ hypothesis test on the thin plate spline fit, using the variance estimates derived in Section 3.5. The issues of performing a pointwise hypothesis test with $N=40000$ points are manifold, and we defer discussion of extensions through which we could make the significance test more sophisticated to Section 12. Nonetheless, we see in Figure 7 a simple fact: that the estimated cooling effect induced by tropical cyclones is of a greater magnitude than the variability of the estimates. Moreover, this cooling effect persists for a sustained period of time, stretching into two weeks following the tropical cyclone passage. We also see a rightward bias in the cooling effect, which may be attributed to the storm’s direction of rotation. [^2] Due to the Coriolis effect, TCs in the Northern Hemisphere will induce a greater cooling effect on the right side of the cross-track angle parameterization than on the left side, due to the combined effects of forward movement and counterclockwise wind rotation. TCs in the Southern Hemisphere, which experience clockwise wind rotation, will induce a greater cooling effect on the left side. In order to display this effect in a consistent manner across hemispheres, we flip the sign of the cross-track angle for TCs in the Southern Hemisphere when producing Figures 612.
@@
This phenomenon may be observed in Figure 8. In Figure 8 we recall the smoothed temperature differences at depth $z=10$, in which a dramatic cooling effect is depicted. At the deeper pressure level of $z=60$, however, we see a mixture of both cooling and warming; namely, a net increase in temperature is observed at $d\in(2,4)$, $\tau\in(0,6)$. Finally, we see in Figure 8 that by $z=150$, the warming effect has disappeared, while the cooling effect remains. We note that the characterization of this warming effect as a continuous function of both cross-track angle and time since TC passage is made possible by our methodological framework and was not possible in previous literature (Cheng et al., 2015).
The discussion of vertical mixing is further developed in Sections 4.3, 4.4, and 4.5. The full set of covariance-reweighed thin plate spline fits, at all pressure levels $z=10,\dotsc,200$, is deferred to Section 2 of the supplementary material (Hu et al., 2020).
@@ 4.4 Depth-cross-track parameterization
We further demonstrate that our continuous-time methodology captures the temporally binned results of Cheng et al. (2015), whose profile pairing process we used in this paper. To introduce temporal binning, we take the covariance-reweighed thin plate spline estimates from Section 4.2 across pressure levels $z=10,\dotsc,200$ and split them into two bins: $\tau\in[0,+3)$ and $\tau\in[+3,+20)$, termed the “forced” stage and “recovery” stage by Cheng et al. (2015). For each of the two bins, we perform the same averaging process as in Section 4.3 to marginalize the estimates over the $\tau$ axis and obtain $\alpha=0.05$ pointwise significance tests.
@@
Several noteworthy phenomena are present in Figure 10. In Figure 10, we observe both the primary cooling effect and the secondary warming effect in the subsurface on the right side of the storm due to vertical mixing, as discussed in Sections 4.2 and 4.3. We also observe a sustained cooling effect in Figure 10, during the period called the “recovery” stage in Cheng et al. (2015). While the primary cooling effect observed in the two time periods is qualitatively consistent with the results presented by Cheng et al. (2015), their results also show a strong warming effect on both sides of the storm during the recovery stage (Figure 10 in Cheng et al., 2015), which is absent from our estimate. Overall, it appears that the recovery stage estimate in Cheng et al. (2015) is “warm biased” in comparison to our estimate. We suspect that this difference is due to our methodology’s more careful accounting of the seasonal mean shift as described in Section 3.3.
@@ 4.5 Isosurfaces of the thin plate spline estimates
((a)) $(\theta,\varphi)=(21.0^{\circ},144.1^{\circ})$
((b)) $(\theta,\varphi)=(25.1^{\circ},16.1^{\circ})$
((c)) $(\theta,\varphi)=(18.6^{\circ},176.4^{\circ})$
@@
For each set of isosurfaces, we first mask the thin plate spline estimates that fail to reject a pointwise level $\alpha=0.05$ hypothesis test, using the variance estimates from Section 3.5.2. We then apply the marching cubes algorithm (Lorensen and Cline, 1987; Lewiner et al., 2012; Virtanen et al., 2020) to identify a positive and a negative isosurface of magnitude $0.3^{\circ}\mathrm{C}$. The positive isosurface is colored in red, and the negative in blue.